OpenStudy (anonymous):

1. Solve 5x2 = –45x 1. Solve 5x2 = –45x @Mathematics

OpenStudy (anonymous):

x^2 = -45x / 5 x = -45/5 x=-9

OpenStudy (anonymous):

when you divide by x, you lose a solution. The original equation is quadratic, so there are two solutions. One of them can be obtained in the manner above, but you lost the other. So solve like this instead, so you get both:\[5x^2=-45x\iff5x^2+45x=0\iff 5x(x+9)=0\iff x=0,x=-9 \]

OpenStudy (anonymous):

either is good answer

OpenStudy (anonymous):

right, both will solve the problem. sometimes the question will want both solutions though. just a heads up :)

OpenStudy (anonymous):

just in case do i add x = 0 to the answer ?

OpenStudy (anonymous):

they are both viable solutions so supplying one or the other wont make you wrong

OpenStudy (anonymous):

ok